OpenStudy (anonymous):

Robert needs enough mulch to cover the triangular garden shown below. If one bag of mulch covers 12 square feet, how many bags of mulch does he need to buy? (Points : 2) 5 11 4 6

OpenStudy (anonymous):

@zordoloom

OpenStudy (anonymous):

@zordoloom

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

What is the area of the triangle? There are 2 of them. Calculate the area of one and then the other.

OpenStudy (anonymous):

12 square feet?

jimthompson5910 (jim_thompson5910):

or you can just calculate the area of that single big triangle A = (b*h)/2 A = (14.1*9)/2 A = 63.45

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

now divide that result by 12 to get 63.45/12 = 5.2875 and round up to get 6

OpenStudy (anonymous):

thts a big number

OpenStudy (anonymous):

but one side is larger than the other.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

31.3 + 31.7 =63. He needs 5.25 bags. I guess @jim_thompson5910 was correct. They are really close.

OpenStudy (anonymous):

6 bags.